# Glonet: Mercator’s end-to-end neural forecasting system

### v1

@@ 4.3.1 Derived quantities
In this part two essential physical properties are derived; geostrophic currents [Rhines, 1979] and mixed layer depth (MLD) [de Boyer Montégut et al., 2004] from the forecasted variables to assess the physical consistency, dynamic accuracy, and potential artifacts in the models outputs.
@@
are derived from forecasted SSH under the geostrophic approximation:
@@
Figure  9 presents examples of derived geostrophic currents computed from GLONET, Xihe, and GLORYS12, the latter serving as a reference. GLONET’s derived geostrophic currents show very clean results, reproducing structures and patterns nearly identical to those in GLORYS12. This indicates a high degree of physical consistency in GLONET’s SSH forecasts, capturing the expected gradient flows with minimal distortion. The similarity in structure between GLONET and GLORYS12 further attests to GLONET’s capability in maintaining dynamic accuracy even when evaluated through derived, process-oriented quantities. In contrast, Xihe’s derived geostrophic currents appear notably noisy, with visible artifacts in the output. These artifacts are largely concealed in direct SSH predictions but become evident when calculating the geostrophic current, a process that involves spatial gradients. Since gradient-based calculations can amplify inconsistencies or abrupt changes, Xihe’s noise suggests underlying instabilities or artifacts in the SSH field that would otherwise remain undetected.
@@
(MLD) is derived from forecasted temperature and salinity profiles and serves as a key indicator of upper-ocean stratification and mixing. MLD is commonly defined based on a density threshold criterion, such that the mixed layer is the depth at which the density difference from the surface equals a specified threshold. The MLD can be approximated as:
@@
In contrast, Xihe’s derived MLD fields are noticeably noisier, with considerable spatial inconsistencies and a lack of coherent structure, indicating underlying artifacts. These artifacts reflect instabilities in Xihe’s forecasted temperature and salinity fields, leading to inconsistencies in the derived density profiles. Such issues underscore the limitations of models like Xihe in capturing the delicate balance between temperature and salinity required for realistic MLD calculations. The observed noise and irregularity in Xihe’s MLD also point to inconsistencies between the forecasted temperature and salinity fields, which can disrupt the physical consistency of derived variables.
@@
In contrast, Xihe displays significantly higher RMSE values for both geostrophic currents and MLD, indicating substantial deviations from GLORYS12. For geostrophic currents, the elevated RMSE suggests notable discrepancies in the SSH-derived fields, potentially due to unresolved artifacts or inconsistencies. Similarly, the high RMSE in MLD forecasts suggests challenges in preserving coherence between temperature and salinity, pointing to physical limitations in Xihe’s derived fields.
@@ 4.4 Diagnostic Variables
Beyond the visual analysis of cascade attenuation, where a progressive loss of small-scale features is observed over time, the energy dynamics is further investigated using a quantitative approach. To assess the retention of small-scale dynamics over time, the energy distribution, derived from the vorticity field, is analyzed across spatial scales. This way the models’ ability to preserve fine-grained dynamics critical for accurate ocean forecasting is evaluated.
The vorticity fields is analyzed in the spectral domain to determine their energy content at different spatial scales. Specifically, the vorticity field $\omega(x,y)$ is subjected to a two-dimensional discrete Fourier transform to obtain its spectral representation $\hat{\omega}(k_{x},k_{y})$, where $k_{x}$ and $k_{y}$ represent the zonal and meridional wavenumbers, respectively. Then the power spectrum is computed as $|\hat{\omega}(k_{x},k_{y})|^{2}$, quantifying the energy associated with each spectral component.
@@
Small-scale energy content is defined as the sum of $E(k)$ over all wavenumbers $k>k_{c}$, where the cutoff wavenumber $k_{c}$ corresponds to a physical length scale $L_{c}$ (e.g., $L_{c}=100\,\text{km}$), and the total energy is computed by summing $E(k)$ across all wavenumbers. The fraction of energy associated with small scales is then defined as:
$E_{\eta}=\frac{\sum\limits_{k>k_{c}}E(k)}{\sum\limits_{k}E(k)}$
This dimensionless fraction quantifies the proportion of total energy contained in small spatial scales relative to the entire spectrum.
By examining the small-scale energy fraction over the 10-days forecast horizon, the ability of each model to maintain the energy cascade characteristic of dynamic fidelity is evaluated. This analysis is conducted for 10-days forecasts spanning the period from January to July 2024, covering both global domains and the Gulf Stream region. The results consistently revealed that both GLONET and GLO12 maintained stable small-scale energy fractions throughout the forecast period, demonstrating their capacity to preserve fine-scale features and sustain energy consistency. Notably, the stability of these fractions was observed across both the global domain and the Gulf Stream region, underscoring the robustness of GLONET and GLO12 in preserving small-scale dynamics.
Conversely, Xihe exhibited a marked decline in the small-scale energy fraction over the forecast horizon, indicative of its inability to sustain the energy cascade. This pattern, consistent across both the global domain and the Gulf Stream region, aligns with the observed decay of vorticity structures in Xihe forecasts. These findings highlight Xihe’s broader challenges in maintaining dynamical coherence and preserving the intricate coupling between scales, essential for accurate ocean forecasting.
@@
The analysis shows that GLONET and GLO12 perform comparably well, accurately capturing the spatial distribution of mesoscale eddies over the 10-day forecast period. Figure  15 illustrates the regional distribution of detected eddies, revealing that both GLONET and GLO12 align closely with the expected patterns of mesoscale variability. This consistency underscores their ability to maintain the SSH gradients necessary for eddy detection and to preserve dynamic integrity across lead times.
@@
Figure 15: Evolution of eddies over a 10-day forecast period for GLONET, GLO12, and Xihe, displayed over the Gulf Stream region. First column represent the eddies extracted from the initial state inputted to all the models.
@@ 4.4.1 Lagrangian analysis
In this subsection, the process-oriented accuracy of GLONET is assessed by analyzing Lagrangian trajectories over a forecast horizon of 10 days, using daily forecast outputs from January to July 2024. This approach offers insight into the model’s ability to capture the advection of ocean particles accurately over time, which is critical for applications that rely on realistic flow and transport simulations.
@@
Figure 16: (Left) Red boxes highlight the areas where particles were initiated. (Right) Divergence of GLONET and Xihe’s Lagrangian trajectories from those of GLORYS12. Trajectories reconstruction is done based on the the weekly forecasts span from from January to July 2024.
@@
The results, illustrated in Figure  17, show that GLONET and GLO12 exhibit consistent RMSE evolution over the 10-day forecast period, with a stable and generally smaller Euclidean distance between geostrophic and full trajectories. This close alignment indicates that both models maintain a high degree of coherence between SSH and current predictions, reflecting an ability to generate forecasts with internally consistent surface and total flow fields. Conversely, Xihe demonstrates a rapid divergence between geostrophic and full current-based trajectories, evident in the increasing Euclidean distance over time. This discrepancy suggests a lack of coherence between Xihe’s SSH and current fields, likely due to artifacts that surface in derived geostrophic currents. The divergence highlights Xihe’s limitations in producing dynamically consistent outputs across interconnected variables.
@@ 5 Discussion and Summary
Data availability statement
GLONET $[1/4^{\circ}]$ experimental daily forecast are accessible through the European Digital Twin Ocean (DTO) platform Edito https://www.edito.eu/.
GLORYS12 and GLO12 products are available in Copernicus Marine Data Store https://data.marine.copernicus.eu/

### v2

@@ 4.3.1 Derived quantities
[…]
@@
provide a diagnostic of large-scale ocean circulation. Accurate prediction of these currents is critical for understanding ocean transport and dynamics. They are derived from forecasted SSH under the geostrophic approximation:
@@
Figure  7 presents examples of derived geostrophic currents computed from GLONET, Xihe, and GLORYS12, the latter serving as a reference. GLONET’s derived geostrophic currents show very clean results, reproducing structures and patterns nearly identical to those in GLORYS12. This indicates a high degree of physical consistency in GLONET’s SSH forecasts, capturing the expected gradient flows with minimal distortion. The similarity in structure between GLONET and GLORYS12 further attests to GLONET’s capability in maintaining dynamic accuracy even when evaluated through derived, process-oriented quantities. In contrast, Xihe’s derived geostrophic currents appear notably noisy, with visible artifacts in the output. These artifacts are largely concealed in direct SSH predictions but become evident when calculating the geostrophic current, a process that involves spatial gradients. Since gradient-based calculations can amplify inconsistencies or abrupt changes, Xihe’s noise suggests underlying instabilities or artifacts in the SSH field that would otherwise remain undetected.
@@
(MLD) is derived from forecasted temperature and salinity profiles and serves as a key indicator of upper-ocean stratification and mixing. Accurately predicting MLD is essential for simulating air-sea interactions, heat exchange, and biological productivity. MLD is commonly defined based on a density threshold criterion, such that the mixed layer is the depth at which the density difference from the surface equals a specified threshold. The MLD can be approximated as:
@@
To ensure a fair and consistent comparison of MLD across all models, we compute the MLD by sampling only the vertical levels that are common to all model configurations within the surface to 600 meters depth range. This approach eliminates discrepancies that could arise from differences in vertical resolution and guarantees that each model’s MLD estimate is based on an identical set of depth points. By standardizing the vertical sampling, we ensure that variations in the derived MLD reflect true model performance rather than artifacts of differing vertical discretizations.
@@
In contrast, Xihe’s derived MLD fields exhibit pronounced noise, characterized by spatial artifacts that predominantly align along latitude and longitude directions. These patterns reflect architectural limitations within Xihe, where instabilities in the forecasted temperature and salinity fields lead to inconsistencies in the reconstructed density profiles. As a result, the MLD fields lack coherent structure and physical realism. The presence of such directional noise highlights challenges in Xihe’s ability to maintain the delicate balance between temperature and salinity, essential for accurate MLD estimation, and points to structural biases introduced by the model’s design.
@@
In contrast, Xihe displays significantly higher RMSE values for the geostrophic currents, and a stable but slightly elevated MLD RMSE, indicating substantial deviations from GLORYS12. For geostrophic currents, the elevated RMSE suggests notable discrepancies in the SSH-derived fields, potentially driven by unresolved artifacts or inconsistencies. In the case of MLD forecasts, the relatively high RMSE is largely attributable to spatial artifacts, primarily aligned along latitude and longitude directions, which disrupt the coherence between temperature and salinity fields. This points to physical limitations in Xihe’s ability to maintain consistency in its derived fields.
@@ 4.4 Diagnostic Variables
Beyond the visual analysis of cascade attenuation, where a progressive loss of small-scale features is observed over time, the energy dynamics is further investigated using a quantitative approach. To assess the retention of small-scale dynamics over time, the energy distribution, derived from the vorticity field, is analyzed across spatial scales. This way, the models’ ability to preserve fine-grained dynamics critical for accurate ocean forecasting is evaluated.
The vorticity fields are analyzed in the spectral domain to determine their energy content at different spatial scales. Specifically, the vorticity field $\omega(x,y)$ is subjected to a two-dimensional discrete Fourier transform to obtain its spectral representation $\hat{\omega}(k_{x},k_{y})$, where $k_{x}$ and $k_{y}$ represent the zonal and meridional wavenumbers, respectively. The power spectrum is then computed as $|\hat{\omega}(k_{x},k_{y})|^{2}$, quantifying the energy associated with each spectral component.
@@
Scale-integrated energy content is defined as the sum of $E(k)$ over all wavenumbers $k<k_{c}$, where the cutoff wavenumber $k_{c}$ corresponds to a physical length scale $L_{c}$ (e.g., $L_{c}=100\,\text{km}$). Note that a cutoff of $L_{c}=100\text{km}$ includes all spectral energy from 100 km and above, while larger cutoff values (e.g., 500 or 1000 km) exclude progressively more small-scale energy. The total energy is computed by summing $E(k)$ across all wavenumbers, and the fraction of energy associated with $L_{c}$ cutoff is then defined as:
$E_{\eta}=\frac{\sum\limits_{k<k_{c}}E(k)}{\sum\limits_{k}E(k)}$ (16)
This dimensionless fraction quantifies the proportion of total energy integrated over spatial scales bounded by $Lc$ relative to the entire spectrum. The goal is not to assess the absolute energy retained but rather whether the spectral shape, particularly at small scales, is preserved across lead times, ideally showing a consistent profile similar to the initial condition.
By examining the scale-integrated energy fraction over the 10 days forecast horizon, the ability of each model to maintain the energy cascade characteristic of dynamic fidelity is evaluated. This analysis is conducted for 10 days forecasts spanning the period from January to July 2024, covering both global domains and the Gulf Stream region. The results consistently revealed that both GLONET and GLO12 maintained stable small-scale energy fractions throughout the forecast period, demonstrating their capacity to preserve fine-scale features and sustain energy consistency. Notably, the stability of these fractions was observed across both the global domain and the Gulf Stream region, underscoring the robustness of GLONET and GLO12 in preserving small-scale dynamics.
Conversely, Xihe exhibited a marked decline in the small-scale energy fraction over the forecast horizon, indicative of its inability to sustain the energy cascade. This pattern, consistent across both the global domain and the Gulf Stream region, aligns with the observed decay of vorticity structures in Xihe forecasts. To further isolate whether this decay is related specifically to the loss of small-scale energy, we repeated the analysis using different cutoff scales ($L_{c}=200$, 500, and 1000 km), as presented in the appendix. Additionally, we conducted a power spectral density (PSD) analysis of the vorticity field to more precisely attribute energy loss to specific spatial scales. These extended analyses confirm that Xihe’s energy decay spans a wide range of scales, highlighting broader challenges in maintaining dynamical coherence and preserving the intricate coupling between scales essential for accurate ocean forecasting.
@@
The analysis shows that GLONET and GLO12 perform comparably well, accurately capturing the spatial distribution of mesoscale eddies over the 10-day forecast period. Figure  13 illustrates the regional distribution of detected eddies, revealing that both GLONET and GLO12 align closely with the expected patterns of mesoscale variability. This consistency underscores their ability to maintain the SSH gradients necessary for eddy detection and to preserve dynamic integrity across lead times.
@@
Figure 13: Evolution of eddies over a 10-day forecast period for GLONET, GLO12, and Xihe, displayed over the Gulf Stream region. The first column represents the eddies extracted from the initial state inputted to all the models.
@@ 4.4.1 Lagrangian analysis
In this subsection, the process-oriented accuracy of GLONET is assessed by analyzing Lagrangian trajectories over a forecast horizon of 10 days, using daily forecast outputs from January to July 2024. Lagrangian drift analysis offers insight into a model’s ability to capture the advection of ocean particles over time, which is critical for applications involving transport processes such as pollutant dispersion, larval connectivity, and passive tracer dynamics. By simulating the motion of synthetic particles advected by model-predicted velocity fields, we assess whether the flow structures are coherent and physically realistic.
@@
Figure 14: (Left) Red boxes highlight the areas where particles were initiated. (Right) Divergence of GLONET and Xihe’s Lagrangian trajectories from those of GLORYS12. Trajectories reconstruction is done based on the weekly forecasts span from from January to July 2024.
@@
The results, illustrated in Figure  15, show that GLONET and GLO12 exhibit consistent RMSE evolution over the 10-day forecast period, with a stable and generally smaller Euclidean distance between geostrophic and full trajectories. This close alignment indicates that both models maintain a high degree of coherence between SSH and current predictions, reflecting an ability to generate forecasts with internally consistent surface and total flow fields. Conversely, Xihe demonstrates a rapid divergence between geostrophic and full current-based trajectories, evident in the increasing Euclidean distance over time. This discrepancy suggests a lack of coherence between Xihe’s SSH and current fields, likely due to artifacts that surface in derived geostrophic currents. The divergence highlights Xihe’s limitations in producing dynamically consistent outputs across interconnected variables.
@@ 5 Discussion and Summary
A key distinction between the three models, GLONET, Xihe, and GLO12 lies in their underlying methodologies. GLONET is a data-driven model trained on historical outputs of physical systems, without direct assimilation of real-time observations or governing physical laws. This leads to its strong performance in large-scale dynamics and ocean currents, but also to a noticeable gap in surface predictions such as SST, where GLO12 benefits from direct assimilation of satellite observations. Xihe, like GLONET, is data-driven, but with a distinct training strategy that uses separate models for each forecast lead time, enabling stable performance across lead times. However, this comes at the cost of dynamic continuity, as demonstrated by Xihe’s weaker performance in preserving derived quantities like geostrophic currents and mesoscale eddies. GLO12, as a physics-based model, excels in surface variables like SST, due to its reliance on both physical equations and real-time data assimilation, making it the most accurate for tropical surface temperature predictions.
@@
Open Research Section
GLONET $[1/4^{\circ}]$ experimental daily forecasts are available on EDITO platform at https://glonet.lab.dive.edito.eu/.
GLORYS12 and GLO12 products are available on the Copernicus Marine Service at https://data.marine.copernicus.eu/.
@@ Acknowledgements
Energy cascade and power spectrum analysis
To complement the main analysis of small-scale energy retention, additional evaluations were performed using alternative cutoff wavelengths ($L_{c}=200$, 500, and 1000 km) to test the sensitivity of the results to spectral thresholding (see Figure 17). These supplementary analyses corroborated the primary findings: both GLONET and GLO12 consistently maintained stable $Lc$-integrated scale energy fractions over the 10-day forecast horizon, indicating a robust capacity to preserve the energy cascade and fine-scale dynamics. Interestingly, at coarser cutoffs (particularly $L_{c}=500$ and $L_{c}=1000$ km), GLONET tends to retain slightly higher energy levels than GLO12, suggesting a marginally more energetic representation at intermediate scales. In contrast, Xihe continued to exhibit a marked decline in energy across all cutoff levels, reinforcing its limitations in sustaining dynamical coherence and multiscale coupling.
Figure 17: Comparison of small-scale energy fraction over the 10-day forecast period for Xihe, GLONET, and GLO12 models. The analysis is based on vorticity fields over the global ocean and spans the evaluation period from January to July 2024. From left to right, panels correspond to cutoff scales of 200 km, 500 km, and 1000 km, respectively.
To further contextualize the small-scale energy retention analysis, we perform a power spectral density (PSD) evaluation of surface vorticity for GLONET, Xihe, and GLO12 at lead times of 1, 5, and 10 days (see Figure 18). This spectral decomposition provides a more detailed view of the energy distribution across spatial scales, offering a complementary perspective to the scale-integrated energy fractions. The results reveal that Xihe systematically loses energy across all wavelengths with increasing lead time, with the depletion being most pronounced at smaller scales (below 200 km). Additionally, Xihe exhibits anomalously high spectral peaks at these finer scales, indicative of noise or unphysical variance accumulating in the absence of dynamical constraints. In contrast, GLO12 maintains a consistent spectral shape across lead times, indicative of stable multiscale dynamics. GLONET displays only mild spectral evolution: while it shows a modest reduction in energy at the smallest scales relative to day-1 forecasts, it retains higher energy at larger scales compared to GLO12. This slight redistribution suggests that GLONET remains energetically coherent over time, albeit with a subtle shift in its spectral energy balance toward broader features.
Figure 18: Power spectral density averaged from January to July 2024 for the entire globe. The analysis is based on vorticity fields and illustrates the distribution of energy across spatial scales at different forecast lead times (day 1, 4, and 9).
Vertical Analysis: GLONET vs GLO12
Building upon the depth-resolved evaluation presented in earlier (Figure 4), a dedicated comparison between GLONET and GLO12, was conducted using GLORYS12 as the reference (see Figure 19). This comparison enables a more balanced assessment of their respective skills across the vertical column. Overall, GLONET demonstrates improved performance in forecasting ocean currents throughout the water column, reflecting its ability to capture and propagate dynamical features more effectively. However, for hydrographic variables such as temperature and salinity, GLONET exhibits slightly lower accuracy than GLO12, particularly at greater depths. These findings suggest that while GLONET offers enhanced dynamical fidelity, it remains more challenged in representing the thermohaline structure of the deep ocean, where GLO12’s physically constrained formulation provides a modest advantage.
Figure 19: RMSE computed at each depth level and averaged over all lead times for 3D variables (U and V currents, temperature, and salinity) for the GLONET (orange line) and GLO12 (green line) models, covering January to July 2024. GLORYS12 serves as the reference, with 10-day forecasts initialized weekly on Wednesdays from a nowcast analysis performed with GLO12 seven days behind real-time.
Architectural Rationale and Component-Level Considerations
Spectral and Spatial Feature Modeling
The spatial architecture of GLONET integrates both spectral and spatial representations to effectively model multi-scale ocean dynamics. The Fourier Neural Operator (FNO) is employed to capture large-scale, low-frequency components by leveraging global receptive fields and efficient spectral parameterizations. While FNO excels at modeling smooth, basin-wide structures such as gyres and planetary waves, its inherent spectral bias limits its capacity to resolve high-frequency, localized phenomena. To mitigate this, a convolutional neural network (CNN) module is incorporated in parallel. The CNN, characterized by localized convolutional kernels, complements the FNO by enhancing sensitivity to sharp gradients and fine-scale structures, including submesoscale eddies and fronts. This dual-branch configuration enables GLONET to capture the full range of spatial variability present in ocean fields.
Temporal Encoding and Implicit Forcing Representation
Temporal information is encoded through a dedicated operator acting on two consecutive states of the ocean. This module is designed to extract the underlying dynamical evolution, thereby implicitly capturing the effects of surface forcing and transient tendencies without the need for explicit atmospheric inputs. By conditioning on both $X_{t-1}$ and $X_{t}$, the temporal encoder learns representations that embed short-term forcings into the model state, which is particularly effective for short-range forecasts where the predictive horizon does not critically depend on long-term external drivers.
Multi-Scale Fusion via Latent-Space Composition
To facilitate integration of representations extracted at disparate spatial scales, GLONET employs an encoder-decoder framework. Rather than aggregating the outputs of the FNO and CNN modules in the physical domain, where representational incompatibilities may impair learning, the model maps both outputs into a shared latent space. This design enables more coherent fusion by aligning features at an abstract representational level, thereby improving the model’s ability to fuse multi-scale information in a physically consistent and data-efficient manner. The decoder then reconstructs the forecasted ocean state from this latent representation, completing the operator mapping from initial condition to future state.
Table 3: List of acronyms.
Acronyms Definition
ACC Antarctic Circumpolar Current
AI Artificial Intelligence
AIFS Artificial Intelligence/Integrated Forecasting System
CNN Convolutional Neural Network
EDITO European Digital Twin Ocean platform EDITO
FNO Fourier neural operators
GDP Global Drifter Program
GLO12 global ocean analysis and forecast system
GLORYS12 Global Ocean Reanalysis and Simulation at $1/12^{\circ}$ resolution (Version 12)
GPU Graphics Processing Unit
HPC High Performance Computing
IV-TT Class IV Task Team
MAE Mean Absolute Error
ML Machine Learning
MLD Mixed Layer Depth
NWP Numerical Weather prediction
RMSE Root Mean Square Error
RMSD Root Mean Square Difference
SLA Sea Level Anomaly
SSH Sea Surface Height
SST Sea Surface Temperature